OpenStudy (anonymous):

A 30- 60- 90 triangle has a hypotenuse of length 9.8. What is the length of the longer of the two legs? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

is length of any other leg given??

OpenStudy (anonymous):

no just the hypotenuse

OpenStudy (anonymous):

cos(30)*9.8

OpenStudy (anonymous):

i got 8.5?

OpenStudy (anonymous):

If C = the hypotenuse, and B = the longer leg, then, sin(60) = B/C = sqrt(3)/2 Substitute C = 9.8: B / 9.8 = sqrt(3)/2 B = (9.8) sqrt(3)/2 = 4.9 sqrt(3) <-----exact answer ~= 8.49 <--------- decimal approximation

OpenStudy (anonymous):

so 8.49 is the required ans

OpenStudy (anonymous):

geometry is your thing huh?

OpenStudy (anonymous):

your a big help!

OpenStudy (anonymous):

well maths in general upto the Indian XIth standard

OpenStudy (anonymous):

u r welcome to all help that i can extend......